An electron moves through a uniform magnetic field given by B = Bxi+(3.74Bx)j. At a particular instant, the electron has velocity V =(2.88i+4.37j)m/s and the magnetic force acting on it is (2.74x10^-19)k N. Find Bx
Answer #1

B=Bxi+3.74Bxj

V=2.88i+4.37j

F=2.74×10-19k N

F=q (V× B)where B and V are vector

V×B=(3.74×2.88k-4.37k)Bx=6.4012Bxk

F=q×6.4012Bxk=1.6×10^-19×6.4012Bxk

Using given value of F we get

2.74×10^-19k=10.24192×10^-19 Bxk

Bx=2.74/10.24192= 0.26753 T